# Improving LLM-based Machine Translation with Systematic Self-Correction

### v2

@@ 5 Analysis
Zero-shot vs Few-shot Prompting
We looked into using both zero-shot and few-shot prompting strategies for our TER framework. The results are shown in Table 1. We noticed that using few-shot prompting strategies generally works better than using zero-shot prompting strategies. Interestingly, we also find that employing zero-shot prompting strategies for estimation can sometimes deteriorate the BLEU score. This suggests, on one hand, the limitations inherent in string-based metrics and, on the other hand, implies potential deficiencies in the quality of zero-shot estimations.
To explore the impact of estimation quality in TER, we utilized GPT-4 with human assistance as gold standard estimation. Figure 5 showcases involving feedback with better estimation quality can help improve the effect of TER in COMET scores. Table 2 shows a comparison between zero-shot and few-shot estimation methods evaluated by human experts. The results favor few-shot approaches, highlighting their superior quality. Additionally, correlation analyses with MQM scores, using GPT-4 with human assistance as gold scores, underscore this finding. We observed that few-shot estimation notably outperforms zero-shot one, as detailed in Table 11. However, the modest correlation scores for few-shot estimation and minor enhancement in TER point to the estimation process as a critical bottleneck in optimizing self-correction efficacy.
@@
Figure 6 present the error analysis with different estimation strategies. We observed that both few-shot GPT-3.5-turbo and GPT-4 with human assistance predominantly identify errors in the categories of Accuracy/Mistranslation and Style/Awkward. The former tends to estimate more errors in the Accuracy/Mistranslation category, while GPT-4 attributes more errors to the Style/Awkward category. Additionally, we find that zero-shot estimation tends to report more errors in categories like Accuracy/Mistranslation, Accuracy/Addition, and Accuracy/Untranslated text, which are considered critical/major in terms of severity. We also provide a typical case in Table 14, which demonstrates the preferences in different estimations. Considering these observations, we infer that weaker estimation strategies tend to overestimate translation errors. Table 3 further illustrates the growth in COMET scores after correcting different types of errors. When error types involve Accuracy, they usually pertain to higher severity levels of errors and be corrected more; whereas error types falling under the Style category are typically associated with lower severity errors.
LLMs Exhibit Diverse Translation and Estimation Capabilities.
Table 4 and 5 present the results about how well different LLMs do in translation and estimation, based on few-shot prompting strategies. We observed that GPT-3.5-turbo performs best in translation for En-De and Zh-En, while Claude-2 excels in En-De and He-En. The average rankings for GPT-3.5-turbo, Gemini-Pro, and Claude-2 are 1.75, 2.5, and 1.75, respectively. We also find that Claude-2 achieves the highest scores in both System-level and Segment-level evaluation (see details in Appendix B) for En-Ru and He-En. As for En-De, the situation is somewhat complex, The ranking order exhibits significant differences between the system-level and segment-level evaluations. We hypothesize that the current MQM is primarily tailored for shorter sentences, potentially leading to reduced robustness when applied to longer paragraph-level tests. We consider both of these two levels in our subsequent analysis.
@@ 6 Related Work
LLMs for Machine Translation
LLM-based MT falls into two main categories. The first focuses on developing strategies, including prompt design, in-context example selection, and more, as outlined in works (Vilar et al., 2023; Agrawal et al., 2023; Zhang et al., 2023; Garcia et al., 2023; Peng et al., 2023b). This category also encompasses the evaluation of MT in various contexts, such as low-resource, document-level, and multilingual translation, as discussed in studies (Hendy et al., 2023; Jiao et al., 2023; Karpinska and Iyyer, 2023; Zhu et al., 2023; Wang et al., 2023). For example, He et al. (2023) employ a selection mechanism based on external quality estimation to filter out unhelpful knowledge. Liang et al. (2023) introduce a Multi-Agent Debate (MAD) framework to encourage divergent thinking in LLMs to reach more accurate translations. The second focuses on training a specific translation LLM. Xu et al. (2023) propose a many-to-many LLM-based translation model, fine-tuning on monolingual data. Yang et al. (2023) construct a comprehensive parallel corpus dataset consisting of 102 languages to enhance the language capabilities of LLaMA. Wu et al. (2024) investigate the adaptation LLMs for document-level machine translation.
@@
Post-hoc Correction
Automatic improvement is a line of research that utilizes LLMs to self-correct outputs. Self-Refine (Madaan et al., 2023) and Reflexion (Shinn et al., 2023) use LLMs to generate output, provide feedback, and refine the output based on that feedback. Recently, Gou et al. (2023) introduce CRITIC, a method designed to improve LLM outputs by integrating external feedback from diverse sources. Huang et al. (2023) examine the ability of LLMs to self-correct their reasoning responses, revealing that LLMs struggle with intrinsic self-correction. In applying post-hoc correction to MT, Chen et al. (2023a) explore using LLMs to iterate the translation, but it lacks clear guidance and the quality improvement is limited. Raunak et al. (2023) compare different post-editing strategies leveraging GPT-4, yet struggle with the lack of precise diagnostics and feedback mechanisms. In contrast, our work introduces a comprehensive self-correction translation framework that clearly separates the estimation module and introduces a structured feedback system, making the refinement process more interpretable and effective.
@@ 7 Conclusion
We introduce a self-correcting translation framework, TER, that employs large language models to perform expert-like guided revisions on translations. Experimental results demonstrate that TER surpasses existing post-editing methods in both metric scores and human preference. We conduct a comprehensive exploration of TER, showcasing the impact of different components on its performance. We delve into the estimate module and present an error-level analysis, demonstrating the key impact of estimation strategies. We also unveil potential connections between the model’s translation capabilities and evaluation proficiency across various language pairs.
@@
Applying self-correction in the field of Machine Translation has inherent advantages, as the problem is well-defined, and the feedback information is highly directional and specific. However, identifying the optimal strategy for applying LLMs to self-correct translations remains a critical area for exploration. As we have uncovered that estimation may become a bottleneck limiting the effectiveness of TER, there is currently no universally proven estimation prompting strategies that demonstrate excellent results. What’s more, due to limited computational resources, we only utilize black-box large language models in this work. It remains uncertain how the TER framework would perform on powerful open-source models.
@@ Appendix A Dataset Statistics
Appendix B Meta Evaluation of MQM
When comparing translation evaluation capabilities of different LLMs, we consider two dimensions used in WMT Shared Metrics task: system-level and segment-level. For the system-level, we utilize pairwise accuracy of system-ranking (Kocmi et al., 2021),
@@ Appendix B Meta Evaluation of MQM
where $\Delta$ represents the difference between the scores of the two systems. For the segment-level, we follow Freitag et al. (2022) to adopt the average of three types of Kendall correlation across all translation pairs.
Appendix C Primary Exploration for Iterative Self-Correction
In our primary experiments, we also explored whether continuously iterating the self-correcting process could enhance performance on the WMT23 Zh-En. We found that this might lead to a decrease in performance, with a single round of self-correcting yielding the best results (see Figure 7). We speculate that low-quality estimation from GPT-3.5 can produce hallucinations, thus leading to poor correction. Additionally, for errors that are easy to identify and correct, one round of self-correction is sufficient. However, for more challenging cases, one round or multiple rounds make a minor difference. We also consider this issue as a direction for future work.
Figure 7: Metric scores for TER as a function of the number of refinement steps on Zh-En. Dashed red line denotes the initial translation baseline.
Appendix D Exploring TER Components from Another Perspective
Table 12 shows the metrics against different parts of initial translations. Zero-shot prompting translations, in comparison to few-shot, have lower scores, indicating poorer quality. Under the same estimation and refinement strategies, the scores for zero-shot IT typically show greater enhancement. We also find that the average score of cases requiring refinement is lower than the all sampled dataset (see Table 1). Additionally, the relative score difference for few-shot is higher than for zero-shot, indicating that a stronger Estimate module (few-shot), compared to a weaker one (zero-shot), can better discriminate and identify cases. Moreover, we observed using feedback alone is sufficient to improve scores, demonstrating the effectiveness of AI feedback in correcting translations. Compared to the feedback-only approach, the $\mathcal{T}_{refine}-\beta$ strategy further enhances the improvement effect.
Appendix E Different LLMs Leading Different Correction Execution Rate
@@
Appendix F Case Study
Table 14 showcases different estimation strategies working on the same initial translation. The weaker model and prompting method make it easy to produce hallucinations and overestimate the translation errors, and also lead to worse downstream refinement.
Appendix G Prompt Templates
Table 15, 16, 17, 18, 19, and 20 show the prompts we used in this work.
Figure 8: Correction execution rate under different initial settings. GPT: GPT-3.5-turbo; GEM: Gemini-Pro; CLA: Claude-2. We first use the same model for translation and estimation. Then, we do self-correction and cross-correction. Succeeded denotes the translation is modified after refinement. Failed represents the refined translation is the same as the initial translation.
@@ Appendix G Prompt Templates
TER 19.37 74.69 76.00 57.21
Table 8: Results of the WMT23 Zh-En refined translations. The best results are in bold. TER is compared with three baseline methods (IT, SCoT, and CT) using GPT-3.5-turbo, Gemini-Pro, and Claude-2.
@@
Table 12: Comparison of quality improvement for different variants of the TER modules on our WMT23 Zh-En refined translations. # of refinements refers to how many cases are judged to need refinement by the Estimate module. $\mathcal{T}_{refine}-\alpha$ uses only the feedback from the Estimate module, $\mathcal{T}_{refine}-\beta$ uses few-shot exemplars in the Translate module and the feedback. $\Delta$ indicates the relative quality against the results without Refine module.
@@
Table 14: Case study of different estimation feedback. Top: translation defects annotated by human expert (in red); Middle: translation defects estimated successfully by models (in red), overestimated errors (under lines); Bottem: unsuccessfully corrected (in red); successfully corrected (in blue). Fdb: source of the estimation feedback.
@@
Please provide the {tgt_lan} translation for the {src_lan} sentences:
Source: {origin}
Target:

### v3

@@ 5 Analysis
Impact of Different TEaR Components
We looked into using both zero-shot and few-shot prompting strategies for each component of our TEaR framework. The results are shown in Table 1. We noticed that using few-shot prompting strategies generally works better than using zero-shot prompting strategies. Interestingly, we also find that employing zero-shot prompting strategies for estimation can sometimes deteriorate the BLEU score. This suggests, on one hand, the limitations inherent in string-based metrics and, on the other hand, implies potential deficiencies in the quality of zero-shot estimations.
To explore the impact of estimation quality in TEaR, we utilized GPT-4 with human assistance as gold standard estimation. Figure 5 showcases involving feedback with better estimation quality can help improve the effect of TEaR in COMET. Table 2 shows human experts favor few-shot estimation approaches, highlighting their superior quality. We also observed that few-shot estimation notably outperforms zero-shot one when calculating the correlation between their MQM scores with the GPT-4 scores, as detailed in Table 11. However, the modest correlation scores for few-shot estimation and minor enhancement in TEaR point to the estimation process as a critical bottleneck in optimizing self-correction efficacy.
@@
As shown in Figure 6, we observed that both few-shot GPT-3.5-turbo and GPT-4 with human assistance predominantly identify errors in the categories of Accuracy/Mistranslation and Style/Awkward. The former tends to estimate more errors in the Accuracy/Mistranslation category, while GPT-4 attributes more errors to the Style/Awkward category. Additionally, we find that zero-shot estimation tends to report more errors in categories like Accuracy/Mistranslation, Accuracy/Addition, and Accuracy/Untranslated text, which are considered as critical/major in terms of severity. We also provide a typical case in Table 19, which demonstrates the preferences in different estimations. Considering these observations, we infer that weaker estimation strategies tend to overestimate translation errors. Appendix E provides more insights to this inference. Table 3 further illustrates the growth in COMET scores after correcting different types of errors. When error types involve Accuracy, they usually pertain to higher severity levels of errors and be corrected more; whereas error types falling under the Style category are typically associated with lower severity errors.
Correlation between Translation and Estimation Capabilities in General-purpose LLMs
Table 4 and 5 present the results about how well different LLMs do in translation and estimation, based on few-shot prompting strategies. We observed that GPT-3.5-turbo performs best in translation for En-De and Zh-En, while Claude-2 excels in En-De and He-En. The average rankings for GPT-3.5-turbo, Gemini-Pro, and Claude-2 are 1.75, 2.5, and 1.75, respectively. We also find that Claude-2 achieves the highest scores in both System-level and Segment-level evaluation (see details in Appendix C) for En-Ru and He-En. As for En-De, the situation is somewhat complex, The ranking order exhibits significant differences between the system-level and segment-level evaluations. We hypothesize that the current MQM is primarily tailored for shorter sentences, potentially leading to reduced robustness when applied to longer paragraph-level tests. We consider both of these two levels in our subsequent analysis.
@@ 6 Related Work
Automatic Post-Editing (APE)
APE aims to correct systematic errors of an MT system whose decoding process is inaccessible and adapt the output to the lexicon and style required in a specific application domain. Traditional APE systems often relied on combining human post-edited data with large, synthetic datasets created through back-translation, a process that is both time-consuming and labor-intensive (Vu and Haffari, 2018; Chatterjee, 2019; Góis et al., 2020). Correia and Martins (2019) utilized a BERT-based model (Devlin et al., 2018) with transfer learning for APE. Chollampatt et al. (2020) explored APE to enhance NMT model translation quality. Several studies (Shterionov et al., 2020; Voita et al., 2019; do Carmo et al., 2020) have examined different neural model architectures and the integration of contextual information to improve post-edited translations.
With the development of LLMs, new learning-based methods have emerged to train LLMs for refining translations (Xu et al., 2023b; Alves et al., 2024; Koneru et al., 2023). Recent research has also explored leveraging powerful LLMs like ChatGPT to refine translations through innovative prompting strategies. Chen et al. (2023) explore using LLMs to iterate translations but lack clear guidance, resulting in limited quality improvement. Raunak et al. (2023) compare different post-editing strategies leveraging GPT-4 but struggle with the lack of precise diagnostics and feedback mechanisms. In contrast, our work introduces a comprehensive self-refinement framework that clearly separates the estimation module and introduces a structured feedback system, making the refinement process more interpretable and effective.
LLMs for Machine Translation (MT)
LLM-based MT falls into two main categories. The first focuses on developing strategies, including prompt design, in-context example selection, and more, as outlined in works (Agrawal et al., 2023; Zhang et al., 2023; Garcia et al., 2023; Peng et al., 2023b; He et al., 2023b; Liang et al., 2023). This category also encompasses the evaluation of MT in various contexts, such as low-resource, document-level, and multilingual translation, as discussed in studies (Hendy et al., 2023; Jiao et al., 2023; Karpinska and Iyyer, 2023; Zhu et al., 2023; Wang et al., 2023). The second focuses on training a specific translation LLM (Zeng et al., 2023; Yang et al., 2023; Xu et al., 2023a; Guo et al., 2024; Alves et al., 2024). For example, Xu et al. (2023a) propose a many-to-many LLM-based translation model, fine-tuning on monolingual data. Wu et al. (2024) investigate the adaptation LLMs for document-level machine translation.
@@ 7 Conclusion
We introduce a self-correcting translation framework, TEaR, that employs large language models to perform expert-like guided revisions on translations. Experimental results demonstrate that TEaR surpasses existing post-editing methods in both metric scores and human preference. We conduct a comprehensive exploration of TEaR, showcasing the impact of different components on its performance. We delve into the estimate module and present an error-level analysis, demonstrating the key impact of estimation strategies. We also unveil potential connections between the model’s translation capabilities and evaluation proficiency across various language pairs.
@@
Applying self-refinement in the field of Machine Translation has inherent advantages, as the problem is well-defined, and the feedback information is highly directional and specific. However, identifying the optimal strategy for applying LLMs to self-correct translations remains a critical area for exploration. As we have uncovered that estimation may become a bottleneck limiting the effectiveness of TEaR, there is currently no universally proven estimation prompting strategies that demonstrate excellent results. What’s more, distilling the performance of TEaR from black-box LLMs to smaller-size LLMs (like 7B or 13B) is also a promising future work.
@@ Appendix A Dataset Statistics
Appendix B Evaluation Methods
Appendix C Meta Evaluation of MQM
We provide MQM human annotation guideline and hierarchy list in Table 20 and 21. When comparing translation evaluation capabilities of different LLMs, we consider two dimensions used in WMT Shared Metrics task: system-level and segment-level. For the system-level, we utilize pairwise accuracy of system-ranking (Kocmi et al., 2021),
@@ Appendix C Meta Evaluation of MQM
where $\Delta$ represents the difference between the scores of the two systems. For the segment level, we follow Freitag et al. (2022) to adopt the average of three types of Kendall correlation across all translation pairs.
Appendix D Primary Exploration for Iterative Self-Correction
In our primary experiments, we also explored whether continuously iterating the self-correcting process could enhance performance on the WMT23 Zh-En. We found that this might lead to a decrease in performance, with a single round of self-correcting yielding the best results (see Table 12). We speculate that low-quality estimation from GPT-3.5 can produce hallucinations, thus leading to poor correction. So we further randomly sampled 50 cases from the WMT23 Zh-En dataset and tested them using GPT-4-0613 (see Table 13). Our results show consistent improvement, indicating that TEaR is effective even with stronger models. We also observed that GPT-4 identified only 7 initial translations as requiring modifications, and it proceeded to make changes to all of them. In an iterative setting, we observed that GPT-4-0613 is more robust compared to GPT-3.5-turbo-0613. That said, an LLM-powered self-correction system with stronger models may be able to detect that the provided translation does not have issues and return back the original translation, rather than hallucinating error annotations and corrections. We consider this issue as a direction for future work.
Appendix E Applying TEaR in Other Models
To observe the performance with TEaR of models with significant differences in capabilities, essentially using TEaR as a function. We selected three models with distinct capabilities: GPT-4-0613, GPT-3.5-turbo-0613, and Mistral-7B (Mistral-7B-Instruct-v0.2). Experiments were conducted on the same dataset as in Appendix D. We found that GPT-4-0613 showed the smallest improvement, while GPT-3.5-turbo-0613 showed the largest (see Table 14). We also analyzed the modification requirements for cases deemed necessary by each LLM during the TEaR process. It was observed that GPT-4-0613 was the model least likely to identify its own translations for correction, while GPT-3.5-turbo-0613 identified the most (see Table 15). We speculate this is because GPT-4-0613 inherently has stronger translation capabilities and produces fewer hallucinations, leading to more cautious estimations. Due to time and resource constraints, we will explore the effects of applying TEaR across models of varying capabilities as our future direction.
Appendix F Exploring TEaR Components from Another Perspective
Table 17 shows the metrics against different parts of initial translations. Zero-shot prompting translations, in comparison to few-shot, have lower scores, indicating poorer quality. Under the same estimation and refinement strategies, the scores for zero-shot IT typically show greater enhancement. We also find that the average score of cases requiring refinement is lower than the all sampled dataset (see Table 1). Additionally, the relative score difference for few-shot is higher than for zero-shot, indicating that a stronger Estimate module (few-shot), compared to a weaker one (zero-shot), can better discriminate and identify cases. Moreover, we observed using feedback alone is sufficient to improve scores, demonstrating the effectiveness of AI feedback in correcting translations. Compared to the feedback-only approach, the $\mathcal{T}_{correct}-\beta$ strategy further enhances the improvement effect.
Appendix G Different LLMs Leading Different Correction Execution Rate
@@
Appendix H Case Study
Table 19 showcases different estimation strategies applied to the same initial translation. Weaker models and prompting methods tend to produce hallucinations and overestimate translation errors, leading to poorer downstream refinement.
Appendix I Prompt Templates
For $\mathcal{T}_{translate}$, we follow the setting in Hendy et al. (2023) and Kocmi et al. (2023), using the few-shot (5-shot) high-quality translation pairs. For $\mathcal{T}_{estimate}$, we use few-shot (3-shot) multi-lingual MQM annotated examples to ask LLMs to estimate the initial target in a reference-free scenario. Kocmi and Federmann (2023) have demonstrated that this prompt method ensures the estimation can be executed across any language pairs and can rival metrics models trained on a large amount of MQM annotated data. Table 22, 23, 24, 25, 26, and 27 show the prompts we used in this work.
Figure 7: Correction execution rate under different initial settings. GPT: GPT-3.5-turbo; GEM: Gemini-Pro; CLA: Claude-2. We first use the same model for translation and estimation. Then, we do self-correction and cross-correction. Modified denotes the translation is modified after refinement. Unmodified represents the refined translation is the same as the initial translation.
@@ Appendix I Prompt Templates
TEaR 19.37 74.69 76.00 57.21
Table 8: Results of the WMT23 Zh-En refined translations. The best results are in bold. TEC is compared with three baseline methods (IT, SCoT, and CT) using GPT-3.5-turbo, Gemini-Pro, and Claude-2.
@@
Metric IT Iter1 Iter2 Iter3 Iter4 Iter5
COMET 78.67 79.50 79.29 79.28 79.12 79.20
COMETKiwi 77.22 78.91 78.42 78.47 78.39 78.35
BLEURT 64.28 65.91 65.65 65.17 65.30 65.01
Table 12: Metric scores for TEaR as a function of the number of refinement steps using GPT-3.5-turbo on our WMT23 Zh-En. IT represents initial translation. Iter represents the iteration step.
BLEURT ($\Delta$BLEURT) IT Iter1 Iter2 Iter3 Iter4 Iter5
GPT-4-0613 70.08 70.20 (+0.12) 70.23 (+0.15) 70.20 (+0.12) 70.23 (+0.15) 70.20 (+0.12)
GPT-3.5-turbo 69.25 69.74 (+0.49) 69.80 (+0.55) 69.60 (+0.35) 69.49 (+0.24) 69.67 (+0.42)
Table 13: Comparison of GPT-4 and GPT-3.5 using iterative refinement on 50 sampled WMT23 Zh-En cases.
Model COMET BLEURT
IT TEaR IT TEaR
GPT-4-0613 82.35 82.38 (+0.03) 70.08 70.20 (+0.12)
GPT-3.5-turbo 81.60 82.02 (+0.42) 69.25 69.74 (+0.49)
Mistral-7B 76.54 76.58 (+0.04) 62.88 63.35 (+0.47)
Table 14: Comparison of different models using TEaR. IT represents initial translation. TEaR refers to the version after refining the initial translation using TEaR.
Model CN CM CU
GPT-4-0613 7 7 0
GPT-3.5-turbo 24 20 4
Mistral-7B 20 19 1
Table 15: The statistics of cases that require refinement after estimation using TEaR in different models. "CN" represents the cases that need to be corrected after estimation, "CM" represents the cases that are modified, and "CU" represents the cases in "CN" but remain unchanged after refinement.
Model IT Win Tie TEaR Win
GPT-4-0613 1 3 3
GPT-3.5-turbo 4 11 9
Mistral-7B 4 9 7
Table 16: Human preference using TEaR in different models.
@@
Table 17: Comparison of quality improvement for different variants of the TEaR modules on our WMT23 Zh-En refined translations. # of refinements refers to how many cases are judged to need refinement by the Estimate module. $\mathcal{T}_{refine}-\alpha$ uses only the feedback from the Estimate module, $\mathcal{T}_{refine}-\beta$ uses few-shot exemplars in the Translate module and the feedback. $\Delta$ indicates the relative quality against the results without Refine module.
@@
Table 19: Case study of different estimation feedback. Top: translation defects annotated by human expert (in red); Middle: translation defects estimated successfully by models (in red), overestimated errors (under lines); Bottom: unsuccessfully corrected (in red); successfully corrected (in blue). Fdb: source of the estimation feedback.
MQM Annotator Guidelines
You will be assessing translations at the segment level, where a segment may contain one or more sentences. Each segment is aligned with a corresponding source segment, and both segments are displayed within their respective documents. Annotate segments in natural order, as if you were reading the document. You may return to revise previous segments.
Please identify all errors within each translated segment, up to a maximum of five. If there are more than five errors, identify only the five most severe. If it is not possible to reliably identify distinct errors because the translation is too badly garbled or is unrelated to the source, then mark a single Non-translation error that spans the entire segment.
To identify an error, highlight the relevant span of text, and select a category/sub-category and severity level from the available options. (The span of text may be in the source segment if the error is a source error or an omission.) When identifying errors, please be as fine-grained as possible. For example, if a sentence contains two words that are each mistranslated, two separate mistranslation errors should be recorded. If a single stretch of text contains multiple errors, you only need to indicate the one that is most severe. If all have the same severity, choose the first matching category listed in the error typology (eg, Accuracy, then Fluency, then Terminology, etc).
Please pay particular attention to document context when annotating. If a translation might be questionable on its own but is fine in the context of the document, it should not be considered erroneous; conversely, if a translation might be acceptable in some context, but not within the current document, it should be marked as wrong.
There are two special error categories: Source error and Non-translation. Source errors should be annotated separately, highlighting the relevant span in the source segment. They do not count against the 5-error limit for target errors, which should be handled in the usual way, whether or not they resulted from a source error. There can be at most one Non-translation error per segment, and it should span the entire segment. No other errors should be identified if Non-Translation is selected.
Table 20: MQM annotator guidelines
Error Category Description
Accuracy Addition Translation includes information not present in the source.
Omission Translation is missing content from the source.
Mistranslation Translation does not accurately represent the source.
Untranslated text Source text has been left untranslated.
Fluency Punctuation Incorrect punctuation (for locale or style).
Spelling Incorrect spelling or capitalization.
Grammar Problems with grammar, other than orthography.
Register Wrong grammatical register (eg, inappropriately informal pronouns).
Inconsistency Internal inconsistency (not related to terminology).
Character encoding Characters are garbled due to incorrect encoding.
Terminology Inappropriate for context Terminology is non-standard or does not fit context.
Inconsistent use Terminology is used inconsistently.
Style Awkward Translation has stylistic problems.
Locale Address format Wrong format for addresses.
convention Currency format Wrong format for currency.
Date format Wrong format for dates.
Name format Wrong format for names.
Telephone format Wrong format for telephone numbers.
Time format Wrong format for time expressions.
Other Any other issues.
Source error An error in the source.
Non-translation Impossible to reliably characterize the 5 most severe errors.
Table 21: MQM hierarchy.
@@
Please provide the {tgt_lan} translation for the {src_lan} sentences: Source: {origin} Target: